# QGFace: Quality-Guided Joint Training For Mixed-Quality Face Recognition

### v1

@@ V Conclusion
We have presented our approach of quality-guided joint training strategy for mixed-quality face recognition. Based on quality partitioning, the high-quality data is supervised with the classification loss while the low-quality data is supervised with the contrastive loss. In this way, we can learn both the HQ and the LQ images simultaneously without any other network structures to cope with the mixed-quality faces in real applications. The proxy-updated real-time contrastive queue avoids the feature whitening problem in the momentum encoder queue and improves the discriminability on the LQ data. Our approach is as elegant as a single feature extractor, and can be easily integrated with other approaches such as super-resolution and distillation to further improve the performance.
TABLE V: Rank-1 accuracies (%) of different methods on Tinyface. * means that the method is trained with ResNet100 on MS1MV2 dataset.
@@
Quality Invariant QualNet50-LM (CVPR21) [47] 35.54
@@
Synthesis CSRI-FT (ACCV18) [23] 44.80
@@
Encoder- Only CurricularFace(CVPR20)* [41] 63.68
@@
Acknowledgments
This work was supported by the National Natural Science Foundation of China (No. 61375016) and ECNU Multifunctional Platform for Innovation (001) Center for High Performance Computing.

### v2

@@ V Conclusion
We have presented our approach of quality-guided joint training strategy for mixed-quality face recognition. Based on quality partitioning, the high-quality data is supervised with the classification loss while the low-quality data is supervised with the contrastive loss. In this way, we can learn both the HQ and the LQ images simultaneously without any other network structures to cope with mixed-quality faces in real applications. The proxy-updated real-time contrastive queue avoids the feature-whitening problem in the momentum encoder queue and improves the discriminability on the LQ data. Our approach is as elegant as a single feature extractor, and it can be easily integrated with other approaches, such as super-resolution and distillation, to further improve the performance.
TABLE V: Rank-1 accuracy of different methods on Tinyface. * means the method is trained with ResNet100 on MS1MV2 dataset.
@@
Quality
Invariant
QualNet50-LM (CVPR21) [16] 35.54
@@
W. Hainan et al. (FG23) [40] 66.3
Z. Meng et al.(TIP24) [49] 54.3
CATFacce* (TBIOM24) [37] 68.95
Synthesis
CSRI-FT (ACCV18) [4] 44.80
@@
Encoder-
Only
CurricularFace* (CVPR20) [12] 63.68